HETAV PATEL

Grade 11,

The mass of caco3 produced when co2 gas is bubbled through 500ml of 0.5M ca(oh)2 will be.(please give short detail also)

The mass of caco3 produced when co2 gas is bubbled through 500ml of 0.5M ca(oh)2 will be.(please give short detail also)

Grade:12

3 Answers

Anil Pannikar AskiitiansExpert-IITB
13 years ago

Dear Sunny,

 

Ca(OH)2 + CO2 = CaCO3 + H2O

0.5 = mass/500*10-3*100

so mass = 25

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Anil Pannikar

IIT Bombay

shreyas ramoji
13 years ago

equivalent wt. of Ca(OH)2 is 36.05 g.

36.05 g of Ca(OH)2  in 1000cm3 (1000 ml) of water gives 1 N solution.

36.05/2 = 18.025 g of Ca(OH)2 in 1000 ml of water gives 0.5 N solution.

so 18.025/2 = 9.0125 g of Ca(OH)2  in 500 ml of water gives 0.5 N solution.

hence, 500 ml solution (0.5 N) contains 9.0125 g of Ca(OH)2.....

 

Ca(OH) + CO2 ----> CaCO3 + H2O....

 

molecular mass of Ca(OH)2  is 72.1

" " of CaCO3 is 100.09

Mass of Ca(OH)2 in 500 ml of water is 9.0125 g

Mass of CaCO3 produced is 9.0125x100.09/72 = 12.517 g (approx).

 

plz approve this answer....




Karthik Eyan
13 years ago

Molarity = no, of moles / Volume of solution [in L]

0.5 M = x moles / 0.5

No. of moles of Ca(OH)2 = 0.25

 

Ca(OH)2 + CO2 --> CaCO3 + H20

According to the balanced stoichiometric equation, 1 mole of Ca(OH)2 gives 1 mole of CaCO3.

0.25 moles of Ca(OH)2 will give 0.25 moles of CaCO3.

0.25 moles of CaCO3 = .25 * 100 = 25 grams

 

Hence, 25 grams of CaCO3 is produced.